OpenStudy (anonymous):

add or subtract the rational expressions, as indicated, and simplify your answer

OpenStudy (anonymous):

Do you have the actual problem?

OpenStudy (anonymous):

yes i just submitted just one moment

OpenStudy (anonymous):

\[\frac{ x+2 }{ x-3? }-\frac{ x-3 }{ x+5? }-\frac{ 1 }{ x^2+2x-15? }\]

OpenStudy (anonymous):

Ok, factor the denominator of the fraction on the right

OpenStudy (anonymous):

ok just one moment..........(x+5) (x-3)

OpenStudy (anonymous):

Perfect. Now you have to make a common denominator in order to cleanly subtract all of these. Any idea on what that can be?

OpenStudy (anonymous):

would the common denominator be the x+5 x-3

OpenStudy (anonymous):

Yep! Now for the first fraction, All you need is an x+5, so you multiply that on the bottom, and you can leave it like (x-3)(x+5). But whatever you do to the bottom, you have to do to the top. So multiply the (x+5) to the top one as well

OpenStudy (anonymous):

so it would look like this....\[\frac{ x+2 }{ x-3? }*\frac{ x+5 }{ ?x+5 }-\frac{ x-3 }{ x+5? }*\frac{ x+3 }{ ?x+3 }\]

OpenStudy (anonymous):

No not really more like this: \[\frac{ (x+2)(x-3) }{ (x-3)(x+5) }- \frac{(x-3)(x+5) }{ (x-3)(x+5) }- \frac{ 1 }{ (x-3)(x+5) }\]

OpenStudy (anonymous):

what did i do wrong can u tell.......im looking at some notes we took in class....did you skip a step?? just asking

OpenStudy (anonymous):

You made the x-3 a x+3, thats all Now just for the first 2 fractions, multiply them out.

OpenStudy (anonymous):

*The numerator only

OpenStudy (anonymous):

so are u saying i multiply the numerator by (x+5)

OpenStudy (anonymous):

The numerators should be multiplied out.. like use FOIL

OpenStudy (anonymous):

ok so x^2-3x+2x-6??

OpenStudy (anonymous):

ok let me try again i think i have it